# Template AASArticle with Examples: v6.3.1111Released on March, 1st, 2021

### v1

@@ Abstract
Many promising explosion models for the elusive origin of Type Ia supernovae (SNe Ia) ultimately fail to completely reproduce a number of observed properties of these events. One limiting factor for many of these models is the use of the local thermodynamic equilibrium (LTE) assumption in the calculation of their synthetic observables, which has been shown to prevent the accurate prediction of a number of fundamental features of SNe Ia. The inclusion of high-accuracy non-LTE physics, however, increases computational cost and complexity such that multidimensional non-LTE calculations are often unfeasible, which can be problematic for models that are inherently multidimensional. In this work, we conduct radiative transfer calculations using 1D profiles that each correspond with a line of sight from an asymmetric, 2D SN Ia model. We find, in LTE, that the synthetic observables from these calculations efficiently reproduce those from the 2D calculation when an equivalence of bolometric luminosities between the 1D and 2D treatments is enforced. This allows for the accurate calculation of synthetic observables in 1D while still preserving multidimensional effects associated with the model. We leverage this to produce high accuracy observables from 1D non-LTE calculations, showing significantly improved agreement with observation, including a roughly 50% reduction of $B$-band decline rate into congruence with the observed Phillips relation. Additionally, our non-LTE observables show Si ii $\lambda$5972 pEWs that are much more similar to observation, while spanning multiple Branch classes, suggesting that some spectral classifications of SNe Ia arise from line of sight effects.
@@
software: Sedona (Kasen et al. 2006), CMFGEN (Hillier & Dessart 2012), Spextractor (github.com/anthonyburrow/spextractor), SNooPy (Burns et al. 2011), matplotlib (Hunter 2007, matplotlib.org).
@@ 1 Introduction
Type Ia supernovae (SNe Ia) are integral to a number of aspects in astronomy and cosmology, including the chemical evolution of the Universe, discovery of the accelerating expansion of the Universe, and determination of the local Hubble constant (Riess et al., 1998; Perlmutter et al., 1999). While it is general consensus that these events arise from the thermonuclear explosion of a carbon/oxygen white dwarf in a binary system, the exact progenitor systems that undergo a SN Ia, as well as the ignition mechanism and type of explosion, remain uncertain. A number of proposed models for SNe Ia have shown promise, however none have yet been able to precisely reproduce the observable characteristics of both individual events and the population of SNe Ia as a whole. A model, or collection of models, that could accurately predict the observed properties of SNe Ia would improve the utility of SNe Ia across numerous areas in astronomy and cosmology.
A wide variety of progenitor systems and explosion models for SNe Ia have been proposed over the last few decades (see Liu et al., 2023, for a recent detailed review). In summary, SNe Ia originate from exploding white dwarfs that may have a mass that is Chandrasekhar or sub-Chandrasekhar and the explosion is triggered via some interaction with its binary companion. In the case of a Chandrasekhar-mass progenitor, the common view is that the central densities increase during mass transfer from a non-degenerate companion until the white dwarf undergoes a deflagration and/or a detonation (Whelan & Iben, 1973; Iben & Tutukov, 1984; Nomoto et al., 1984; Hoeflich et al., 1995; Röpke et al., 2007; Seitenzahl et al., 2012; Sim et al., 2013). Alternatively for sub-Chandrasekhar progenitors, the currently favored scenario is the double detonation, in which the explosion of the white dwarf is triggered by the detonation of its helium-rich shell, which is initiated during mass transfer from its likely degenerate companion (Nomoto, 1982; Woosley et al., 1986; Fink et al., 2010; Polin et al., 2019; Shen et al., 2021b; Collins et al., 2022; Shen et al., 2024). It is plausible that both of these progenitor channels are responsible for observed SNe Ia (Maoz et al., 2014).
Many of these proposed models show potential in their synthetic observables with moderate agreement with some individual observed events. However, broad reproduction of the properties of SNe Ia as a whole remains elusive. For example, while the double detonation scenario has proven an ability to generate spectra that emulate individual events across a range of peak magnitudes, the Phillips and color-magnitude relations of observed SNe Ia are not yet well-replicated by multidimensional models (Shen et al., 2021b; Collins et al., 2022).
@@
This has borne out in some recent works that have shown distinct differences in the synthetic observables between LTE and non-LTE calculations of sub-Chandrasekhar detonation models. Shen et al. (2021a) showed a notable difference in the spectra and light curves from 1D sub-Chandrasekhar explosion models between LTE and non-LTE calculations beyond maximum light, especially within $U$- and $B$-band. Blondin et al. (2022) also showed in a radiative transfer code comparison study that CMFGEN (Hillier & Dessart, 2012), a non-LTE code, produces synthetic observables from a sub-Chandrasekhar SN Ia ejecta profile that have generally faster rises to peak and bluer colors compared to most LTE codes. Additionally, Collins et al. 2023 identified a non-LTE-exclusive feature, He I $\lambda 10830$, which may be a unique signature of the double detonation scenario. We refer the reader to Dessart et al. (2014) for a detailed exploration into the critical ingredients for non-LTE calculations of a SN Ia model.
@@
Given the described disparity in observed properties of SNe Ia and what is obtainable from present day calculations of theoretical models, we seek to develop a method in which non-LTE physics is taken into consideration in radiative transfer calculations of a multidimensional SN Ia model, allowing for the generation of accurate synthetic observables from such a model.
In this work, we describe a new method that strives to produce high-accuracy, non-LTE observables from a multidimensional SN Ia model using relatively modest computational resources. In short, this method involves a combination of 1D LTE, 2D LTE, and 1D non-LTE radiative transfer calculations to produce “pseudo-2D” non-LTE synthetic observables. We describe our one examined model and radiative transfer methods in Section 2. In Section 3, we describe the synthetics observables that arise from our calculations and compare them to observed SNe Ia in Section 4. Finally, in Section 5, we discuss our conclusions from this work.
@@ 2 Methods
We describe our choice of SN Ia model for this study in Section 2.1. We also show how we compute the LTE and non-LTE synthetic observables in Sections 2.2 and 2.3, respectively.
@@ 2.1 Double Detonation Model
For this initial work, we probe a single 2D double detonation model from Boos et al. (2021). We choose the 1.0 $M_{\odot}$ total mass WD with a 0.016 $M_{\odot}$ helium shell. This model has a shell base density of $2\times 10^{5}$ g cm-3, which is the lowest shell density examined in Boos et al. 2021. While the helium shell in the double detonation has historically been thought to necessitate a build up in mass from accretion prior to ignition, Shen et al. (2024) recently showed that the natal shell with a realistic composition profile of a degenerate progenitor up to $\sim 1.0$ M is capable of hosting a detonation, thus we choose the double detonation model with the lowest shell mass available.
Figure 1: 2D density and composition ejecta maps for the thin shell, 1.0 $M_{\odot}$ total mass WD model from Boos et al. (2021). The orange lines indicate the boundaries of the 15 wedges that are used to construct the 1D profiles for our 1D calculations, which are numbered in orange in the first frame.
The final ejecta structure in velocity space is displayed in Figure 1, showing density and a selection of notable elements and isotopes. The core ashes are bounded by the innermost layer of carbon, which originates from the outermost layer of the carbon/oxygen core which is not completely burned. The structure of the core ejecta is what can be expected from most sub-Chandrasekhar detonation models, where the innermost region is comprised of 56Ni and other iron group elements. Outward of this region, the ejecta transitions to intermediate mass elements like silicon. Beyond the core ejecta are the products of the helium detonation, which are dependent on its initial density, with more massive shells producing non-trivial amounts of iron group elements. The model selected for this work had a fairly thin shell with a peak density of 2×$\times 10^{5}$ g cm-3 at the core-shell interface and total mass of 0.016 M. Thus, the shell ashes are predominately comprised of intermediate mass elements around the core-shell boundary and unburned helium at higher velocities. This ejecta is a relatively standard expectation for the thin-shelled double detonation scenario, including the stratified core and shell regions as well as their asymmetry, where the velocities are higher in the positive z direction, towards the original helium shell ignition point (Fink et al., 2010; Tanikawa et al., 2018; Gronow et al., 2021; Ferrand et al., 2022).
@@ 2.2 LTE Calculations
The time dependent 2D LTE calculation for our model of focus from Boos et al. (2021) appears in Shen et al. (2021b) using Sedona (Kasen et al., 2006). In addition to that 2D calculation, 1D LTE calculations are carried out here in steady state using Sedona for each of the 15 wedge profiles generated from the 2D ejecta (areas between adjacent pairs of orange lines in Figure 1). For each wedge, we generate spectra every 1.5 days between three days before and 15 days after the time of maximum $B$-band magnitude from the corresponding line of sight in the 2D calculation.
We inject additional power at the origin of the steady-state 1D calculations such that the outgoing bolometric luminosity is within 0.5% of that from the corresponding line of sight in the 2D calculation at the same time. This added power is primarily a stand-in for the deposited energy from radioactive decays prior to calculation time, which is otherwise accounted for in the 2D time-dependent calculation. The value is also somewhat dependent on how similar a given wedge is relative to nearby wedges. For example, the northernmost wedge has the lowest amount of 56Ni and necessitates relatively more injected energy in the 1D calculations due to higher amounts of 56Ni in nearby wedges that otherwise increase the luminosity at the corresponding line of sight in the 2D calculation. The amount of additional power needed to match the 2D calculation is less at later times due to the larger contribution of the instantaneous radioactive decay luminosity to the outgoing luminosity and decreasing overall opacity of the ejecta. The power is injected via a blackbody of temperature $3\times 10^{4}$ K, but we find relatively little effect from this choice on the synthetic observables. Our imposed constraint on the bolometric luminosity is supported by previous works that show fairly consistent bolometric light curves between LTE and non-LTE calculations of SNe Ia models within a month of the explosion (Dessart et al., 2014; Shen et al., 2021a; Blondin et al., 2022; Sharon & Kushnir, 2024), allowing for bolometric unification between our 1D and 2D calculations.
@@ 2.3 Non-LTE Calculations
In this section we describe the procedure used for evolving the ejecta structures along each wedge of the 2D explosion model with the radiative transfer code CMFGEN  (Hillier & Miller, 1998; Dessart & Hillier, 2005; Hillier & Dessart, 2012). We used the standard time-dependent non-LTE solver in CMFGEN with an approach and setup analogous to those used in previous Type Ia SN studies (see, e.g., Blondin et al. 2017a). A few days before $B$-band maximum, the same wedge structure that is used in Sedona is imported into CMFGEN and remapped onto an optical-depth scale counting about 90 depth points and covering from 250 until about 45,000 km s-1. For the composition, we account for the presence of He, C, N, O, Ne, Na, Mg, Al, Si, S, Ar, Ca, Sc, Ti, V, Cr, Mn, Fe, Co, and Ni. We consider three two-step decay chains with parent isotopes 56Ni, 48Cr, and 52Fe. Non-thermal rates arising from Compton-scattered, high-energy electrons are computed through a solution to the Spencer-Fano equation, as discussed in Li et al. (2012) and Dessart et al. (2012).
@@
We employ an extended model atom with the following ions, super-levels and full levels (see Hillier & Miller 1998 for details about the treatment of full and super levels): He i (40,51), He ii (13,30), C i (14,26), C ii (14,26), C iii (62,112), N i (44,104), N ii (23,41), N iii (25,53), O i (19,51), O ii (30,111), O iii (50,86), Ne i (70,139), Ne ii (22,91), Ne iii (23,71), Na i (22,71), Mg ii (22,65), Mg iii (31,99), Al ii (26,44), Al iii (17,45), Si ii (31,59), Si iii (33,61), Si iv (37,48), S ii (56,324), S iii (48,98), S iv (27,67), Ar i (56,110), Ar ii (134,415), Ar iii (32,346), Ca ii (21,77), Ca iii (16,40), Ca iv (18,69), Sc ii (38,85), Sc iii (25,45), Ti ii (37,152), Ti iii (33,206), V i (1,1), Cr ii (28,196), Cr iii (30,145), Cr iv (29,234), Mn ii (25,97), Mn iii (30,175), Fe i (44,136), Fe ii (275,827), Fe iii (83,698), Fe iv (100,1000), Fev (47,191), Co ii (136,2747), Co iii (124,3917), Co iv (37,314), Cov (32,387), Ni ii (59,1000), Ni iii (47,1000), Ni iv (36,200), and Niv (46,183).
@@
We start the CMFGEN simulations for each wedge at 10–15 d after explosion and typically a few days before $B$-band maximum for that given wedge (as computed by Sedona). For the first time step in the sequence, we compute the model with the temperature fixed and assuming steady state. Once this initial model is converged, we evolve the model with time dependence using a time increment equal to 10 % of the current time until we reach about 30-35 d. The goal is to cover from $B$-band maximum until 15 d later in the CMFGEN sequence (the time of $B$-band maximum computed by Sedona and CMFGEN differ).
@@
For each wedge and time, we initially compute the $\gamma$-ray energy deposition profile using a pure-absorption, grey transport solver with $\gamma$-ray opacity equal to 0.06 $Y_{\rm e}$ cm2 g-1, where $Y_{\rm e}$ is the material electron fraction. Once this model is converged, we rerun it by scaling the depth-dependent energy-deposition profile so that the volume integrated decay power equals the 2D luminosity computed by Sedona for that wedge direction. With this approach, we mimic in 1D the luminosity of the 2D model while capturing the critical non-LTE effects inherent to the radiative transfer.

### v2

# Non-LTE Synthetic Observables of a Multidimensional Model of Type Ia Supernovae

@@ Abstract
Many promising explosion models for the elusive origin of Type Ia supernovae (SNe Ia) ultimately fail to completely reproduce a number of observed properties of these events. One limiting factor for many of these models is the use of the local thermodynamic equilibrium (LTE) assumption in the calculation of their synthetic observables, which has been shown to prevent the accurate prediction of a number of fundamental features of SNe Ia. The inclusion of high-accuracy non-LTE physics, however, increases computational cost and complexity such that multidimensional non-LTE calculations are often unfeasible, which can be problematic for models that are inherently multidimensional. In this work, we conduct radiative transfer calculations using 1D profiles that each correspond with a line of sight from an asymmetric, 2D SN Ia model. We find, in LTE, that the synthetic observables from these calculations efficiently reproduce those from the 2D calculation when an equivalence of bolometric luminosities between the 1D and 2D treatments is enforced. This allows for the accurate calculation of synthetic observables in 1D while still preserving multidimensional effects associated with the model. We leverage this to produce high accuracy observables from 1D non-LTE calculations, showing significantly improved agreement with observation, including a roughly 50% reduction of $B$-band decline rate into congruence with the observed Phillips relation. Additionally, our non-LTE observables show Si ii $\lambda$5972 pEWs that are much more similar to observation, while spanning multiple Branch classes, suggesting that some spectral classifications of SNe Ia may arise from line of sight effects.
@@
software: Sedona (D. Kasen et al. 2006), CMFGEN (D. J. Hillier & L. Dessart 2012), Spextractor (github.com/anthonyburrow/spextractor), SNooPy (C. R. Burns et al. 2011), matplotlib (J. D. Hunter 2007, matplotlib.org).
@@ 1 Introduction
Type Ia supernovae (SNe Ia) are integral to a number of aspects in astronomy and cosmology, including the chemical evolution of the Universe, discovery of the accelerating expansion of the Universe, and determination of the local Hubble constant (A. G. Riess et al., 1998; S. Perlmutter et al., 1999). While it is general consensus that these events arise from the thermonuclear explosion of a carbon/oxygen white dwarf in a binary system, the exact progenitor systems that undergo a SN Ia, as well as the ignition mechanism and type of explosion, remain uncertain. A number of proposed models for SNe Ia have shown promise, however none have yet been able to precisely reproduce the observable characteristics of both individual events and the population of SNe Ia as a whole. A model, or collection of models, that could accurately predict the observed properties of SNe Ia would improve the utility of SNe Ia across numerous areas in astronomy and cosmology.
A wide variety of progenitor systems and explosion models for SNe Ia have been proposed over the last few decades (see Z.-W. Liu et al., 2023, for a recent detailed review). In summary, SNe Ia originate from exploding white dwarfs that may have a mass that is Chandrasekhar or sub-Chandrasekhar and the explosion is triggered via some interaction with its binary companion. In the case of a Chandrasekhar-mass progenitor, the common view is that the central densities increase during mass transfer from a non-degenerate companion until the white dwarf undergoes a deflagration and/or a detonation (J. Whelan & J. Iben, 1973; J. Iben & A. V. Tutukov, 1984; K. Nomoto et al., 1984; P. Hoeflich et al., 1995; F. K. Röpke et al., 2007; I. R. Seitenzahl et al., 2012; S. A. Sim et al., 2013). Alternatively for sub-Chandrasekhar progenitors, a favored scenario is the double detonation, in which the explosion of the white dwarf is triggered by the detonation of its helium-rich shell, which is initiated during mass transfer from its likely degenerate companion (K. Nomoto, 1982; S. E. Woosley et al., 1986; M. Fink et al., 2010; A. Polin et al., 2019; K. J. Shen et al., 2021b; C. E. Collins et al., 2022; K. J. Shen et al., 2024). It is plausible that both of these progenitor channels are responsible for observed SNe Ia (D. Maoz et al., 2014).
Many of these proposed models show potential in their synthetic observables with moderate agreement with some individual observed events. However, broad reproduction of the properties of SNe Ia as a whole remains elusive. For example, while the double detonation scenario has proven an ability to generate spectra that emulate individual events across a range of peak magnitudes, the Phillips and color-magnitude relations of observed SNe Ia are not yet well-replicated by multidimensional models (K. J. Shen et al., 2021b; C. E. Collins et al., 2022).
@@
This has borne out in some recent works that have shown distinct differences in the synthetic observables between LTE and non-LTE calculations of sub-Chandrasekhar detonation models. K. J. Shen et al. (2021a) showed a notable difference in the spectra and light curves from 1D sub-Chandrasekhar explosion models between LTE and non-LTE calculations beyond maximum light, especially within $U$- and $B$-band. S. Blondin et al. (2022) also showed in a radiative transfer code comparison study that CMFGEN (D. J. Hillier & L. Dessart, 2012), a non-LTE code, produces synthetic observables from a sub-Chandrasekhar SN Ia ejecta profile that have generally faster rises to peak and bluer colors compared to most LTE codes. Additionally, C. E. Collins et al. 2023 identified a non-LTE-exclusive feature, He I $\lambda 10830$, which may be a unique signature of the double detonation scenario. We refer the reader to L. Dessart et al. (2014a) for a detailed exploration into the critical ingredients for non-LTE calculations of a SN Ia model.
@@
Given the described disparity in observed properties of SNe Ia and what is obtainable from present day calculations of theoretical models, we seek to develop a method in which non-LTE physics is taken into consideration in radiative transfer calculations of a multidimensional SN Ia model, allowing for the generation of accurate synthetic observables from such a model. In this work, we describe a new method that strives to produce high-accuracy, non-LTE observables from a multidimensional SN Ia model using relatively modest computational resources. In short, this method involves a combination of 1D LTE, 2D LTE, and 1D non-LTE radiative transfer calculations to produce “pseudo-2D” non-LTE synthetic observables. We describe our one examined model and radiative transfer methods in Section 2. In Section 3, we describe the synthetics observables that arise from our calculations and compare them to observed SNe Ia in Section 4. Finally, in Section 5, we discuss our conclusions from this work.
@@ 2 Methods
We describe our choice of SN Ia model for this study in Section 2.1. We also show how we compute the LTE and non-LTE synthetic observables in Sections 2.2 and 2.3, respectively. Finally, we describe our strategy to construct pseudo-2D non-LTE observables in Section 2.4.
@@ 2.1 Double Detonation Model
For this initial work, we probe a single 2D double detonation model from S. J. Boos et al. (2021). We choose the 1.0 $M_{\odot}$ total mass WD with a 0.016 $M_{\odot}$ helium shell. This model has a shell base density of $2\times 10^{5}$ g cm-3, which is the lowest shell density examined in S. J. Boos et al. 2021. While the helium shell in the double detonation has historically been thought to necessitate a build up in mass from accretion prior to ignition, K. J. Shen et al. (2024) recently showed that the natal shell with a realistic composition profile of a degenerate progenitor up to $\sim 1.0$ M is capable of hosting a detonation, thus we choose the double detonation model with the lowest shell mass available.
Figure 1: 2D density and composition ejecta maps at approximately 100 s post-explosion for the thin shell, 1.0 $M_{\odot}$ total mass WD model from S. J. Boos et al. (2021). The orange lines indicate the boundaries of the 15 wedges that are used to construct the 1D profiles for our 1D calculations, which are numbered in orange in the first frame.
The final ejecta structure in velocity space is displayed in Figure 1, showing density and a selection of notable elements and isotopes. The core ashes are bounded by the innermost layer of carbon, which originates from the outermost layer of the carbon/oxygen core which is not completely burned. The structure of the core ejecta is what can be expected from most sub-Chandrasekhar detonation models, where the innermost region is comprised of 56Ni and other iron group elements. Outward of this region, the ejecta transitions to intermediate mass elements like silicon. Beyond the core ejecta are the products of the helium detonation, which are dependent on its initial density, with more massive shells producing non-trivial amounts of iron group elements. The model selected for this work had a fairly thin shell with a peak density of 2×$\times 10^{5}$ g cm-3 at the core-shell interface and total mass of 0.016 M. Thus, the shell ashes are predominately comprised of intermediate mass elements around the core-shell boundary and unburned helium at higher velocities. This ejecta is a relatively standard expectation for the thin-shelled double detonation scenario, including the stratified core and shell regions as well as their asymmetry, where the velocities are higher in the positive z direction, towards the original helium shell ignition point (M. Fink et al., 2010; A. Tanikawa et al., 2018; S. Gronow et al., 2021; G. Ferrand et al., 2022).
@@ 2.2 LTE Calculations
The time dependent 2D LTE calculation for our model of focus from S. J. Boos et al. (2021) appears in K. J. Shen et al. (2021b) using Sedona (D. Kasen et al., 2006). In addition to that 2D calculation, 1D LTE calculations are carried out here using Sedona for each of the 15 wedge profiles constructed from the 2D ejecta (areas between adjacent pairs of orange lines in Figure 1). For each wedge, we generate spectra every 1.5 days between three days before and 15 days after the time of maximum $B$-band magnitude from the corresponding line of sight in the 2D calculation.
A typical 1D radiative transfer calculation using these wedge profiles will generally misestimate the total outgoing luminosity compared to the 2D result at the corresponding line of sight. This is due to the fact that contributions from adjacent wedges affect the outgoing luminosity for a given line of sight in the 2D calculation. For example, the northernmost wedge has the lowest amount of 56Ni and will significantly underestimate the outgoing luminosity in the 1D steady-state treatment due to higher amounts of 56Ni in nearby wedges that increase the luminosity at the corresponding line of sight in the 2D calculation. Thus, we elect to conduct these 1D LTE radiative transfer calculations in steady-state, allowing us to modulate the outgoing luminosity at each time step. Aside from energy contributions from adjacent wedges, a typical steady-state calculation in the photospheric phase will also generally underestimate the total outgoing luminosity due to a lack of consideration of the deposited energy from radioactive decays prior to the calculation time, which is otherwise accounted for in a time-dependent calculation.
To unify the 1D LTE luminosities with corresponding 2D LTE values, we increase the power of our 1D LTE calculations by scaling the average radioactive decay energy in the $\rm{{}^{56}Ni-^{56}Co-^{56}Fe}$ decay chain. We modulate this scaling factor for each wedge and timestep using a Newton-Raphson iteration, until the outgoing bolometric luminosity of each 1D steady-state calculation is within 0.5% of that from the corresponding line of sight in the 2D calculation at the same time. The amount of additional power needed to match the 2D calculation is less at later times due to the larger contribution of the instantaneous radioactive decay luminosity to the outgoing luminosity and decreasing overall opacity of the ejecta. We show an alternative, but less robust, method to unify the 1D steady-state and 2D time-dependent luminosities in Appendix C. Our imposed constraint on the bolometric luminosity is supported by previous works that show fairly consistent bolometric light curves between LTE and non-LTE calculations of SNe Ia models within a month of the explosion (L. Dessart et al., 2014a; K. J. Shen et al., 2021a; S. Blondin et al., 2022; A. Sharon & D. Kushnir, 2024), allowing for bolometric unification between our 1D and 2D calculations. To avoid confusion with our truly 2D LTE results, we refer to these calculations and resulting synthetic observables as “1D” LTE, even though they are enhanced in such a way to ideally replicate a 2D result.
@@ 2.3 Non-LTE Calculations
In this section we describe the procedure used for evolving the ejecta structures along each wedge of the 2D explosion model with the radiative transfer code CMFGEN  (D. J. Hillier & D. L. Miller, 1998; L. Dessart & D. J. Hillier, 2005; D. J. Hillier & L. Dessart, 2012). We used the standard time-dependent non-LTE solver in CMFGEN with an approach and setup analogous to those used in previous Type Ia SN studies (see, e.g., S. Blondin et al. 2017). A few days before $B$-band maximum, the same wedge structure that is used in Sedona is imported into CMFGEN and remapped onto an optical-depth scale counting about 90 depth points and covering from 250 until about 45,000 km s-1. For the composition, we account for the presence of He, C, N, O, Ne, Na, Mg, Al, Si, S, Ar, Ca, Sc, Ti, V, Cr, Mn, Fe, Co, and Ni. We consider three two-step decay chains with parent isotopes 56Ni, 48Cr, and 52Fe. Non-thermal rates arising from Compton-scattered, high-energy electrons are computed through a solution to the Spencer-Fano equation, as discussed in C. Li et al. (2012) and L. Dessart et al. (2012).
@@
We employ an extended model atom with the following ions, super-levels and full levels (see D. J. Hillier & D. L. Miller 1998 for details about the treatment of full and super levels): He i (40,51), He ii (13,30), C i (14,26), C ii (14,26), C iii (62,112), N i (44,104), N ii (23,41), N iii (25,53), O i (19,51), O ii (30,111), O iii (50,86), Ne i (70,139), Ne ii (22,91), Ne iii (23,71), Na i (22,71), Mg ii (22,65), Mg iii (31,99), Al ii (26,44), Al iii (17,45), Si ii (31,59), Si iii (33,61), Si iv (37,48), S ii (56,324), S iii (48,98), S iv (27,67), Ar i (56,110), Ar ii (134,415), Ar iii (32,346), Ca ii (21,77), Ca iii (16,40), Ca iv (18,69), Sc ii (38,85), Sc iii (25,45), Ti ii (37,152), Ti iii (33,206), V i (1,1), Cr ii (28,196), Cr iii (30,145), Cr iv (29,234), Mn ii (25,97), Mn iii (30,175), Fe i (44,136), Fe ii (275,827), Fe iii (83,698), Fe iv (100,1000), Fev (47,191), Co ii (136,2747), Co iii (124,3917), Co iv (37,314), Cov (32,387), Ni ii (59,1000), Ni iii (47,1000), Ni iv (36,200), and Niv (46,183). We note that the atomic data for our CMFGEN calculations differs from our Sedona calculations.
@@
We start the CMFGEN simulations for each wedge at 10–15 d after explosion and typically a few days before $B$-band maximum for that given wedge (as computed by Sedona). For the first time step in the sequence, we compute the model with the temperature fixed and assuming steady-state. Once this initial model is converged, we evolve the model with time dependence using a time increment equal to 10 % of the current time until we reach about 30-35 d. The goal is to cover from $B$-band maximum until 15 d later in the CMFGEN sequence (the time of $B$-band maximum computed by Sedona and CMFGEN differ).
@@
For each wedge and time, we initially compute the $\gamma$-ray energy deposition profile using a pure-absorption, grey transport solver with $\gamma$-ray opacity equal to 0.06 $Y_{\rm e}$ cm2 g-1, where $Y_{\rm e}$ is the material electron fraction. Once this model is converged, we rerun it by scaling the depth-dependent energy-deposition profile so that the volume integrated decay power (i.e. outgoing luminosity) equals the 2D outgoing luminosity computed by Sedona for that wedge direction. With this approach, we mimic in 1D the luminosity of the 2D model while capturing the critical non-LTE effects inherent to the radiative transfer. Similar to our 1D LTE results, we likewise refer to these calculations and observables as “1D” non-LTE, despite their multidimensional considerations.
2.4 Construction of Pseudo-2D non-LTE Observables
In this work, we aim to simultaneously analyze non-LTE and multidimensional contributions to the synthetic observables. As proper multidimensional, non-LTE calculations are generally unfeasible at this time, we concatenate our three sets of radiative transfer results to produce an approximated facsimile of such a calculation. We refer to this new set of synthetic observables as “pseudo-2D non-LTE”. Specifically, we take the differences in flux at each wavelength between the 1D and 2D LTE calculations (the red bands in Figure 2, i.e. the multidimensional correction) and apply them to the 1D non-LTE spectra at each time step for each line of sight. Likewise, for the theoretical photometry of our models, we take the flux difference between the 1D LTE and 2D LTE in each band and apply them to the 1D non-LTE to construct our pseudo-2D non-LTE light curves. The total energy of these pseudo-2D non-LTE observables are conserved under this strategy (i.e. the bolometric light curves between all four sets of synthetic observables are equivalent within a negligible tolerance). The 2D LTE calculation has a time step of 0.5 d, while the 1D LTE and non-LTE calculations have a time step of 1.5 d, in sync with every third 2D LTE time.
The utility of the pseudo-2D non-LTE synthetic observables is marginally limited in that their construction is not physically motivated. Additionally, there may be non-LTE processes that are exclusively relevant in multiple dimensions. However, these corrected observables are useful for identifying regions of spectra and times where multidimensional features may not be captured by the 1D non-LTE calculations. We include these pseudo-2D non-LTE synthetic observables in comparisons with observed SNe Ia spectra, light curves, and population correlations in following sections.